# Unconditional Density Bounds for Quadratic Norm-Form Energies via Lorentzian Spectral Weights Persistent Heuristics: Mini Monograph I

### 03 Mar 2026 v2

@@ 8.3. The containment lemma and density bound
The weights $b_{k}=w(\gamma_{k}^{\prime})$, the tail $\epsilon_{M}=\sum_{k>M}b_{k}$, and the density $\|f_{S_{L}^{(M)}}\|_{\infty}$ all depend on $d$ through the zeros of $L(s,\chi_{d})$. Theorem 8.5 is valid for every $d$ and every $M$, but the numerical evaluations in this paper ($\epsilon_{20}$, $f_{S_{L}^{(20)}}(0)=8.3129$, and $C(5)=0.1191$) are computed for $d=5$. At fixed $M=20$, the constant $B_{20}$ exceeds $1$ for $d=5$ (since $\epsilon_{20}=0.0687$), making the unconditional bound numerically vacuous at this truncation level. The bound becomes nontrivial for $d\gtrsim 10^{4}$ at $M=20$. For the $O(1/\sqrt{d})$ rate (Corollary 8.8), $M$ must grow with $d$, with the required growth rate depending on the zero distribution of $L(s,\chi_{d})$.
@@
Under the hypothesis that the infinite resonance lattice $\Lambda_{\infty}$ has finite rank $d_{\infty}<\infty$ (in fact $d_{\infty}=0$, verified for $M\leq 20$), then
@@
The density bound relies on the zero set $\{S_{L}=0\}$, which has codimension $1$ and can accumulate mass under resonances. By contrast, the zero set $\{Z_{L}=0\}=\{S_{L}=0\}\cap\{\widetilde{S}_{L}=0\}$ has codimension $2$ in the torus. Crucially, a resonance relation $\sum n_{k}\gamma_{k}^{\prime}=0$ constrains the cosine sum via $\cos(\sum n_{k}\gamma_{k}^{\prime}\delta)=1$ but constrains the sine sum via $\sin(\sum n_{k}\gamma_{k}^{\prime}\delta)=0$; these are algebraically distinct constraints. For instance, the relation $\gamma_{1}^{\prime}=2\gamma_{2}^{\prime}$ forces $\cos(\gamma_{1}^{\prime}\delta)=2\cos^{2}(\gamma_{2}^{\prime}\delta)-1$, but the corresponding sine identity $\sin(\gamma_{1}^{\prime}\delta)=2\sin(\gamma_{2}^{\prime}\delta)\cos(\gamma_{2}^{\prime}\delta)$ has a different zero set. As a consequence, $S_{L}$ and $\widetilde{S}_{L}$ cannot both vanish simultaneously at generic points of the orbit, even in the presence of resonances: the codimension-$2$ structure ensures $K_{L}(\delta)>0$ generically.
@@ 9.1. Verified resonance absence
No exact relation was found. The conditions of Proposition 9.1 are therefore satisfied at $M_{1}=M_{2}=20$.
@@ 9.2. The exact density formula
With asymptotic independence verified at $M=20$, the marginal densities $f_{S_{\zeta}}$ and $f_{S_{L}}$ are both computable via inverse Fourier transform:
@@
$\operatorname{dens}\{N_{d}>0\}=\frac{C}{\sqrt{d}}+o(1/\sqrt{d}),$
where $C$ is computable from the exact marginal densities:
$C=\lim_{d\to\infty}\sqrt{d}\cdot 2\int_{0}^{W_{1}(L)}f_{S_{L}}(y)\,\overline{F}_{|S_{\zeta}|}(\sqrt{d}\,y)\,dy,$ (13)
where $\overline{F}_{|S_{\zeta}|}(u)=P(|S_{\zeta}|>u)$ denotes the survival function of $|S_{\zeta}|$.
@@
which identifies $C=2f_{S_{L}}(0)\int_{0}^{W_{1}(\zeta)}\overline{F}_{|S_{\zeta}|}(u)\,du$.
The factored form $C(d)=2f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|]$ (Corollary 9.5), with both factors computed by rigorous ARB quadrature, gives $C(5)=0.1191$ certified to four significant figures. ∎
@@
$C=2\,f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|],$ (14)
@@
Numerically, $f_{S_{L}}(0)=8.3129$ and $\mathbb{E}[|S_{\zeta}|]=0.00716$, giving $C(5)=0.1191$. Both values were computed for $d=5$ via rigorous ARB quadrature (acb_calc_integrate) at 256-bit precision; see Appendix A, §A.6. These are the exact $M\to\infty$ limiting values, not truncation-level approximations: the telescoping bound (§7) certifies that corrections from zeros beyond $M=20$ contribute less than $10^{-3}$ to $f_{S_{L}}(0)$ and less than $10^{-5}$ to $C$.
@@
The proof of Theorem 9.4 identifies $C=2f_{S_{L}}(0)\int_{0}^{W_{1}(\zeta)}\overline{F}_{|S_{\zeta}|}(u)\,du$. Since $\overline{F}_{|S_{\zeta}|}(u)=0$ for $u>W_{1}(\zeta)$, the upper limit may be replaced by $\infty$. For any non-negative Cesàro-distributed quantity $X$ with finite mean, $\mathbb{E}[X]=\int_{0}^{\infty}\overline{F}_{X}(u)\,du$. Applying this to $X=|S_{\zeta}|$ gives (14).
@@
The factored form (14) clarifies the relationship between the three formulas for the leading constant:
@@
Unconditional bound $2f_{S_{L}}(0)\cdot W_{1}(\zeta)=0.531$ $\sup|S_{\zeta}|$ replaces $\mathbb{E}[|S_{\zeta}|]$
@@
Exact $2f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|]=0.1191$ ($d=5$) (none)
The unconditional bound exceeds the exact value by a factor of $W_{1}(\zeta)/\mathbb{E}[|S_{\zeta}|]=4.46$. This is the reciprocal of the containment tightness ratio $\mathbb{E}[|S_{\zeta}|]/W_{1}(\zeta)=0.224$ (Remark 8.10). The entire slack in the unconditional bound therefore arises from the containment step, which replaces $\mathbb{E}[|S_{\zeta}|]$ with $\sup|S_{\zeta}|$. The Gaussian approximation overestimates $C$ by $22\%$ (predicting $0.145$ instead of $0.1191$). The factored form decomposes this discrepancy: the Gaussian density $1/(\sqrt{2\pi}\,\sigma_{L})$ overestimates $f_{S_{L}}(0)$ by $27\%$ (compact support flattens the Jessen–Wintner density at the origin), while the Gaussian mean $\sigma_{\zeta}\sqrt{2/\pi}$ underestimates $\mathbb{E}[|S_{\zeta}|]$ by $4\%$ (compact support concentrates mass). The two errors partially cancel.
@@
Two numerical values for $f_{S_{L}}(0)$ appear in this paper and refer to the same quantity at different stages of the argument. The value $8.3129$ (Proposition 6.3 and §A.7) is the signed Bessel product integral $(1/\pi)\int_{0}^{T}\prod_{k=1}^{20}J_{0}(b_{k}t)\,dt$ at finite truncation $M=20$, computed by rigorous ARB quadrature to 44 decimal places. The value $8.3129$ (Corollary 9.5) is the density of the limiting Jessen–Wintner distribution at the origin; the $M=20$ truncation captures this to the displayed precision, consistent with the convergence from below established in the telescoping argument. Only the value $8.3129$ enters the formula $C(5)=2f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|]=0.1191$.
@@ 10. Numerical Verification
All numerical results in this section are derived from the ARB-certified ingredients established in the preceding sections: the Besicovitch constants $\sigma_{\zeta}$, $\sigma_{L}$ (Section 4), the unsigned Bessel product integrals $f_{\mathrm{main}}^{(M)}$6), the tail sums $\epsilon_{M}$7), and the leading constant $C(5)=2f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|]=0.1191$ (Corollary 9.5).
@@ 10.1. Density verification
The three columns display successively tighter bounds: the energy ratio (Theorem 8.3), the density bound $2\,f_{\mathrm{main}}^{(20)}\cdot W_{1}(\zeta)/\sqrt{d}$ (Theorem 8.5 at $M=20$), and the exact leading term $C/\sqrt{d}$ (Theorem 9.4).
@@ 11. Conclusion
Stage 1: Finite bounds. At each truncation level $M$, the Jacobi–Anger decomposition combined with Bessel decay estimates and lattice counting proves $f_{S_{L}^{(M)}}(0)<\infty$ unconditionally (Theorem 6.9). The inactive $J_{0}$ factors provide exponential suppression at rate $\eta=e/4\approx 0.68$.
Stage 2: Finite truncation and stability. At any fixed verified truncation level $M$, the effective density bound (Theorem 8.5) gives $\operatorname{dens}\{N_{d}>0\}\leq A_{M}/\sqrt{d}+B_{M}$, which is nontrivial for sufficiently large $d$. The $O(1/\sqrt{d})$ rate (Corollary 8.8) follows under the hypothesis that $\Lambda_{\infty}$ has finite rank $d_{\infty}<\infty$ (verified $d_{\infty}=0$ for $M\leq 20$), which provides a uniform bound on the density and allows the truncation level to grow with $d$.
Stage 3: Verified refinements. Under computationally verified resonance absence at $M=20$, the telescoping argument (Theorem 7.7) establishes $f_{S_{L}}(0)<\infty$ for the full infinite series, and the exact formula $\operatorname{dens}\{N_{d}>0\}=C(d)/\sqrt{d}+o(1/\sqrt{d})$ follows from the computable marginal densities (Theorem 9.4), with $C(5)=0.1191$. No result in this paper assumes GRH or the Grand Simplicity Hypothesis.
@@ 12.2. The unique null crossing
Values of $L(1,\chi_{d})$ were computed from the class number formula $L(1,\chi_{d})=2hR/\sqrt{|\Delta_{K}|}$, where $h$ is the class number and $R=\log\varepsilon$ is the regulator with fundamental unit $\varepsilon$ of $\mathcal{O}_{K}$; all five fields have $h=1$. The null crossings $s_{*}(d)$ were found by bisecting $G(s)=\zeta(s)/L(s,\chi_{d})$ against the target $\sqrt{d}$, with $L(s,\chi_{d})$ evaluated via the Hurwitz decomposition $L(s,\chi_{d})=q^{-s}\sum_{a=1}^{q}\chi_{d}(a)\,\zeta(s,a/q)$ using ARB interval arithmetic at 512-bit precision ($200$ bisection steps, residual $|G(s_{*})-\sqrt{d}|<10^{-60}$).
@@ 12.3. Integer crossing obstruction
Case 2: $s=2$ and $d\neq 3$. The Klingen–Siegel theorem gives $L(2,\chi_{d})=c_{d}\cdot\pi^{2}/\sqrt{d}$ for an explicit rational $c_{d}$ depending on $d$. The null crossing condition $\zeta(2)^{2}=d\cdot L(2,\chi_{d})^{2}$ becomes $\pi^{4}/36=c_{d}^{2}\cdot\pi^{4}$, requiring $c_{d}=1/6$. This holds only for $d=3$.

### 12 Mar 2026 v3

@@ 8.3. The containment lemma and density bound
The weights $b_{k}=w(\gamma_{k}^{\prime})$, the tail $\epsilon_{M}=\sum_{k>M}b_{k}$, and the density $\|f_{S_{L}^{(M)}}\|_{\infty}$ all depend on $d$ through the zeros of $L(s,\chi_{d})$. Theorem 8.5 is valid for every $d$ and every $M$, but the numerical evaluations in this paper ($\epsilon_{20}$, $f_{S_{L}^{(20)}}(0)=8.3129$, and $C(5)=0.1193$) are computed for $d=5$. At fixed $M=20$, the certified value $\epsilon_{20}=0.0308$ gives $B_{20}=2\|f_{S_{L}^{(20)}}\|_{\infty}\cdot\epsilon_{20}\leq 2\times 10.517\times 0.0308\approx 0.648<1$, so the unconditional bound is already nontrivial (strictly below $1$) at $M=20$ for all $d\geq 1$. For the $O(1/\sqrt{d})$ rate (Corollary 8.8), $M$ must grow with $d$, with the required growth rate depending on the zero distribution of $L(s,\chi_{d})$.
@@
Under the hypothesis that the infinite resonance lattice $\Lambda_{\infty}$ has finite rank $d_{\infty}<\infty$ (computationally consistent with $d_{M}=0$ for $M\leq 20$; the hypothesis asserts this extends to the full infinite lattice), then
@@
The density bound relies on the zero set $\{S_{L}=0\}$, which has codimension $1$ and can accumulate mass under resonances. By contrast, the zero set $\{Z_{L}=0\}=\{S_{L}=0\}\cap\{\widetilde{S}_{L}=0\}$ has codimension $2$ in the torus. A resonance relation $\sum n_{k}\gamma_{k}^{\prime}=0$ constrains the cosine sum via $\cos(\sum n_{k}\gamma_{k}^{\prime}\delta)=1$ but constrains the sine sum via $\sin(\sum n_{k}\gamma_{k}^{\prime}\delta)=0$; these are algebraically distinct constraints. For instance, the relation $\gamma_{1}^{\prime}=2\gamma_{2}^{\prime}$ forces $\cos(\gamma_{1}^{\prime}\delta)=2\cos^{2}(\gamma_{2}^{\prime}\delta)-1$, but the corresponding sine identity $\sin(\gamma_{1}^{\prime}\delta)=2\sin(\gamma_{2}^{\prime}\delta)\cos(\gamma_{2}^{\prime}\delta)$ has a different zero set. As a consequence, $S_{L}$ and $\widetilde{S}_{L}$ cannot both vanish simultaneously at generic points of the orbit, even in the presence of resonances: the codimension-$2$ structure ensures $K_{L}(\delta)>0$ generically.
@@ 9.1. Verified resonance absence
No exact relation was found; every combination was certified nonzero by ARB interval arithmetic at 256-bit precision. The conditions of Proposition 9.1 are therefore satisfied at $M_{1}=M_{2}=20$.
@@ 9.2. The exact density formula
Under the hypothesis that the conditions of Proposition 9.1 hold for all $M_{1},M_{2}\geq 1$ (which Proposition 9.3 requires for passing to infinite sums, and which has been verified computationally at truncation level $M=20$), the marginal densities $f_{S_{\zeta}}$ and $f_{S_{L}}$ are both computable via inverse Fourier transform:
@@
$\operatorname{dens}\{N_{d}>0\}=\frac{C(d)}{\sqrt{d}}+o(1/\sqrt{d}),$
where $C(d)$ depends on $d$ (through $f_{S_{L}}(0)$, which depends on the spectral weights $b_{k}=2/({\tfrac{1}{4}}+\gamma_{k}^{\prime 2})$) and is computable from the exact marginal densities:
$C(d)=\lim_{d\to\infty}\sqrt{d}\cdot 2\int_{0}^{W_{1}(L)}f_{S_{L}}(y)\,\overline{F}_{|S_{\zeta}|}(\sqrt{d}\,y)\,dy,$ (13)
where $\overline{F}_{|S_{\zeta}|}(u)=P(|S_{\zeta}|>u)$ denotes the survival function of $|S_{\zeta}|$. For $d=5$, $C(5)=0.1193$ (Section 12).
@@
which identifies $C(d)=2f_{S_{L}}(0)\int_{0}^{W_{1}(\zeta)}\overline{F}_{|S_{\zeta}|}(u)\,du$.
The factored form $C(d)=2f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|]$ (Corollary 9.5), with both factors computed by rigorous ARB quadrature, gives $C(5)=0.1193$ certified to four significant figures. ∎
@@
$C(d)=2\,f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|],$ (14)
@@
Numerically, $f_{S_{L}}(0)=8.3129$ and $\mathbb{E}[|S_{\zeta}|]=0.00717$, giving $C(5)=0.1193$. Both values were computed for $d=5$ via rigorous ARB quadrature (acb.integral) at 256-bit precision; see Appendix A, §A.6. These are the exact $M\to\infty$ limiting values, not truncation-level approximations: the telescoping bound (§7) certifies that corrections from zeros beyond $M=20$ contribute less than $10^{-3}$ to $f_{S_{L}}(0)$ and less than $10^{-5}$ to $C$.
@@
The proof of Theorem 9.4 identifies $C(d)=2f_{S_{L}}(0)\int_{0}^{W_{1}(\zeta)}\overline{F}_{|S_{\zeta}|}(u)\,du$. Since $\overline{F}_{|S_{\zeta}|}(u)=0$ for $u>W_{1}(\zeta)$, the upper limit may be replaced by $\infty$. For any non-negative Cesàro-distributed quantity $X$ with finite mean, $\mathbb{E}[X]=\int_{0}^{\infty}\overline{F}_{X}(u)\,du$. Applying this to $X=|S_{\zeta}|$ gives (14).
@@
The factored form (14) clarifies the relationship between the three formulas for the leading constant (all evaluated at truncation level $M=20$ for both $S_{\zeta}$ and $S_{L}$):
@@
Unconditional bound $2f_{S_{L}}(0)\cdot W_{1}^{(20)}(\zeta)=0.531$ $\sup|S_{\zeta}^{(20)}|$ replaces $\mathbb{E}[|S_{\zeta}^{(20)}|]$
@@
Exact $2f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|]=0.1193$ ($d=5$) (none)
The unconditional bound exceeds the exact value by a factor of $W_{1}^{(20)}(\zeta)/\mathbb{E}[|S_{\zeta}^{(20)}|]=4.45$. This is the reciprocal of the containment tightness ratio $\mathbb{E}[|S_{\zeta}^{(20)}|]/W_{1}^{(20)}(\zeta)=0.225$ (Remark 8.10). The entire slack in the unconditional bound therefore arises from the containment step, which replaces $\mathbb{E}[|S_{\zeta}|]$ with $\sup|S_{\zeta}|$. The Gaussian approximation overestimates $C$ by $22\%$ (predicting $0.145$ instead of $0.1193$). The factored form decomposes this discrepancy: the Gaussian density $1/(\sqrt{2\pi}\,\sigma_{L})$ overestimates $f_{S_{L}}(0)$ by $27\%$ (compact support flattens the Jessen–Wintner density at the origin), while the Gaussian mean $\sigma_{\zeta}\sqrt{2/\pi}$ underestimates $\mathbb{E}[|S_{\zeta}|]$ by $5\%$ (compact support concentrates mass). The two errors partially cancel.
@@
Two numerical values for $f_{S_{L}}(0)$ appear in this paper and refer to the same quantity at different stages of the argument. The value $8.3129$ (Proposition 6.3 and §A.7) is the signed Bessel product integral $(1/\pi)\int_{0}^{T}\prod_{k=1}^{20}J_{0}(b_{k}t)\,dt$ at finite truncation $M=20$, computed by rigorous ARB quadrature to 44 decimal places. The value $8.3129$ (Corollary 9.5) is the density of the limiting Jessen–Wintner distribution at the origin; the $M=20$ truncation captures this to the displayed precision, consistent with the convergence from below established in the telescoping argument. Only the value $8.3129$ enters the formula $C(5)=2f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|]=0.1193$.
@@ 10. Numerical Verification
All numerical results in this section are derived from the ARB-certified ingredients established in the preceding sections: the Besicovitch constants $\sigma_{\zeta}$, $\sigma_{L}$ (Section 4), the unsigned Bessel product integrals $f_{\mathrm{main}}^{(M)}$6), the tail sums $\epsilon_{M}$7), and the leading constant $C(5)=2f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|]=0.1193$ (Corollary 9.5).
@@ 10.1. Density verification
The three columns display successively tighter bounds: the energy ratio (Theorem 8.3), the density bound $2\,f_{\mathrm{main}}^{(20)}\cdot W_{1}^{(20)}(\zeta)/\sqrt{d}$ (Theorem 8.5 at $M=20$, with $W_{1}^{(20)}(\zeta)=\sum_{k=1}^{20}a_{k}=0.03192$), and the exact leading term $C/\sqrt{d}$ (Theorem 9.4).
@@ 11. Conclusion
Stage 1: Finite bounds. At each truncation level $M$, the Jacobi–Anger decomposition combined with Bessel decay estimates and lattice counting proves $f_{S_{L}^{(M)}}(0)<\infty$ unconditionally (Theorem 6.9). The inactive $J_{0}$ factors provide the polynomial $t$-decay needed for integrability, while the subcritical bound on active Bessel factors yields exponential suppression in resonance order at rate $\eta=e/4\approx 0.68$.
Stage 2: Finite truncation and stability. At any fixed verified truncation level $M$, the effective density bound (Theorem 8.5) gives $\operatorname{dens}\{N_{d}>0\}\leq A_{M}/\sqrt{d}+B_{M}$, which is nontrivial for sufficiently large $d$. The $O(1/\sqrt{d})$ rate (Corollary 8.8) follows under the hypothesis that $\Lambda_{\infty}$ has finite rank $d_{\infty}<\infty$ (a hypothesis extending the finite truncation fact $d_{M}=0$, verified for $M\leq 20$), which provides a uniform bound on the density and allows the truncation level to grow with $d$.
Stage 3: Verified refinements. Under computationally verified resonance absence at $M=20$, the telescoping argument (Theorem 7.7) establishes $f_{S_{L}}(0)<\infty$ for the full infinite series, and the exact formula $\operatorname{dens}\{N_{d}>0\}=C(d)/\sqrt{d}+o(1/\sqrt{d})$ follows from the computable marginal densities (Theorem 9.4), with $C(5)=0.1193$. No result in this paper assumes GRH or the Grand Simplicity Hypothesis.
@@ 12.2. The unique null crossing
Values of $L(1,\chi_{d})$ were computed via the digamma formula $L(1,\chi_{d})=-\tfrac{1}{q}\sum_{a=1}^{q}\chi_{d}(a)\,\psi(a/q)$, where $\psi$ denotes the digamma function; each value $\psi(a/q)$ is evaluated as a certified ARB ball via acb.digamma. The null crossings $s_{*}(d)$ were found by bisecting $G(s)=\zeta(s)/L(s,\chi_{d})$ against the target $\sqrt{d}$, with $L(s,\chi_{d})$ evaluated via the Hurwitz decomposition $L(s,\chi_{d})=q^{-s}\sum_{a=1}^{q}\chi_{d}(a)\,\zeta(s,a/q)$ using ARB interval arithmetic at 512-bit precision ($220$ bisection steps from brackets of width 0.50.51.51.5 surrounding each known approximate location, residual $|G(s_{*})-\sqrt{d}|<10^{-60}$). The full certification, including bracket verification, decisive trichotomy bisection, denominator safety via interval union, and the direct integer-crossing check for $d=3$, is provided in the ancillary scripts.
@@ 12.3. Integer crossing obstruction
Case 2: $s=2$ and $d\neq 3$. The Klingen–Siegel theorem gives $L(2,\chi_{d})=c_{d}\cdot\pi^{2}/\sqrt{d}$ for an explicit rational $c_{d}$ depending on $d$. The null crossing condition $\zeta(2)^{2}=d\cdot L(2,\chi_{d})^{2}$ becomes $\pi^{4}/36=c_{d}^{2}\cdot\pi^{4}$, requiring $c_{d}=1/6$.
For all squarefree $d\geq 22$, this is impossible. Since $\chi_{d}(1)=1$ and $|\chi_{d}(n)|\leq 1$ for all $n$, separating the $n=1$ term gives $L(2,\chi_{d})\geq 1-\sum_{n=2}^{\infty}n^{-2}=2-\pi^{2}/6$ for every primitive real character. Therefore $c_{d}\geq(2-\pi^{2}/6)\sqrt{d}/\pi^{2}$, and $c_{d}=1/6$ would require $\sqrt{d}\leq\pi^{2}/(12-\pi^{2})$. That $(2-\pi^{2}/6)\sqrt{22}/\pi^{2}>1/6$ is certified by ARB: the left side equals $0.16874\ldots$ against $1/6=0.16666\ldots$, with certified margin $0.00207$ and ball radius $4.54\times 10^{-154}$. Hence $c_{d}>1/6$ for all squarefree $d\geq 22$.
For the twelve remaining squarefree values $d\in\{2,5,6,7,10,11,13,14,15,17,19,21\}$, the value $c_{d}=\sqrt{d}\,L(2,\chi_{d})/\pi^{2}$ is computed via the Hurwitz decomposition
$L(2,\chi_{d})=q^{-2}\sum_{a=1}^{q}\chi_{d}(a)\,\zeta(2,a/q)$
using ARB interval arithmetic at 512-bit precision, and the ARB ball for $c_{d}$ is certified disjoint from $1/6$, i.e., either $c_{d}>1/6$ or $c_{d}<1/6$ is certified as an ARB predicate. The Klingen–Siegel theorem implies that $c_{d}$ is in fact a rational number for each $d$; the values obtained numerically are listed in the table below for the reader’s convenience, but the certified proof requires only disjointness from $1/6$:
$d$ $\Delta_{K}$ $c_{d}$ $d$ $\Delta_{K}$ $c_{d}$
22 88 $1/8$ $13$ $13$ $4/13$
55 55 $4/25$ $14$ $56$ $5/14$
66 $24$ $1/4$ $15$ $60$ $2/5$
77 $28$ $2/7$ $17$ $17$ $8/17$
$10$ $40$ $7/20$ 1919 $76$ $1/2$
$11$ $44$ $7/22$ 2121 2121 $8/21$
None of these equals $1/6$, completing the case.